a force of 7 newton act on the body the displacement is 8m in the direction of force calculate work done
 

Grade:9

2 Answers

Arun
4 years ago
Dear Ayush
 
F = 7 N
s = 8 m

        W = F × s

        = 7 N × 8 m

        W = 56J

        ∴ Work done is 56 J
 
Hope it helps
 
In case of any difficulty, Please feel free to ask.
Vikas TU
4 years ago
Dear student 
Work done = FS Cos Q 
Q = 0 here 
so work done = 8*7 = 56 J ….................
Hope this helps 
Good Luck